OpenStudy (anonymous):

f(x)=sin(3x)cos(x/3) , find f'(x) , show steps .

OpenStudy (anonymous):

use the product rule f(x)=uv f'(x)= u'v + uv' here u=sin(3x) u'=3cos(3x) v=cos(x/3) v'=-1/3 sin (x/3) f'(x) = 3cos(3x)*cos(x/3) +1/3 sin(3x)(-sin(x/3)

OpenStudy (anonymous):

Thanks!